# High resolution quantum enhanced phase imaging of cells

### v1

@@ I Introduction
To understand why this is the case, it is essential to consider how intensity noise propagates into the retrieved phase profile. By examining Eq. (2), under the assumption that the finite-difference term $I(\mathbf{x},dz)-I(\mathbf{x},-dz)$ is dominated by noise fluctuations $\sigma(\mathbf{x})$, and considering a nearly uniform illumination intensity $I(\mathbf{x},0)\approx I_{0}$, the equation simplifies to Paganin et al. (2004):
$-k\frac{\sigma(\mathbf{x})}{\sqrt{2}\,I_{0}\,dz}=\nabla^{2}\phi_{\text{noise}}(\mathbf{x}).$ (5)
@@
where $\tilde{\sigma}(\mathbf{q})$ and $\tilde{\phi}_{\text{noise}}(\mathbf{q})$ denote the Fourier transforms of the intensity noise and the resulting phase noise, respectively, and $\mathbf{q}$ represents the spatial frequency vector. In the case of shot-noise we can assume a flat spectrum, $\tilde{\sigma}(\mathbf{q})=\tilde{\sigma}$, for $\mathbf{q}\neq 0$. The presence of the $|\mathbf{q}|^{2}$ term in the denominator highlights that the phase noise in the TIE reconstruction is dominated by low spatial frequencies, i.e. by the slowly varying shot-noise fluctuations across the image. In contrast, high-frequency noise components, such as pixel-to-pixel shot-noise, are strongly suppressed by the intrinsic low-pass filtering nature of the TIE, and therefore have minimal impact on the retrieved phase. Incidentally, the remaining low-frequency components of the shot-noise are precisely those efficiently removed through quantum correlations, by the substitution in Eq. (4).
@@
It is noteworthy to mention that the intrinsic suppression of high-frequency noise does not prevent the retrieval of high-resolution phase images. In fact, as it is clear from Eq. (2), in the direct-propagation problem, higher-frequency components of the phase object produce a stronger effect on the intensity. Thus, regions with rapid phase variations, i.e. high-frequency features, are usually reconstructed accurately despite the suppression introduced by solving the TIE Paganin et al. (2004).
Figure 3: Phase resolution evaluation with the optimized correction factor $k_{\text{TIE}}$ (a) Experimental and simulated phase resolution as a function of the resolution factor $\mathcal{D}$ for different defocus distances $dz$. The points correspond to experimental data for various $dz$ values, while shaded areas represent theoretical predictions. As $\mathcal{D}$ increases, the resolution worsens due to the larger effective pixel size, while smaller $dz$ values allow better resolution. (b) Representative phase reconstructions for two different defocus distances ($dz=0.025$ mm and $dz=0.1$ mm) and three different $\mathcal{D}$ values. Insets show magnified regions of the sample for a better qualitative evaluation of the phase resolution changes.
These results are shown in Fig. 1c. The quantum advantage is quantified as the ratio between two Pearson correlation coefficients: one calculated using quantum-corrected phase images and the other using classical (shot-noise limited) phase images, each compared with a reference obtained by averaging $10^{3}$ frames to eliminate residual shot-noise. The Pearson correlation coefficient $\mathcal{C}$, a widely used metric to evaluate the similarity between two images, is defined as:
$\mathcal{C}=\frac{\sum_{x}\left(\phi_{r}(x)-\overline{\phi}_{r}\right)\left(\phi(x)-\overline{\phi}\right)}{\sqrt{\text{Var}[\phi_{r}]\,\text{Var}[\phi]}},$ (7)
@@
In the figure, the blue data points represent the experimental quantum advantage obtained using the optimization factor $k^{(\text{TIE})}_{\text{opt}}=\eta_{0}$ in Eq. (4), with a single-channel overall detection efficiency estimated as $\eta_{0}\approx 0.7$, while the red points indicate results obtained with $k^{(\tau)}_{\text{opt}}=\eta_{0}\eta_{c}(\mathcal{D})$, used for transmittance measurements and previously employed for phase imaging in Ref. Ortolano et al. (2023b). The shaded red and blue areas represent results from Fourier-optics simulations Voelz (2011), showing good agreement with the experimental data, whereas the green area indicates the gain in the quantum advantage achievable with the new reconstruction strategy.
@@
The data are displayed as a function of the defocusing distance $dz$ and the resolution factor $\mathcal{D}$. We observe that using $k^{(\tau)}_{\text{opt}}(\mathcal{D})$, originally defined for amplitude imaging, in the context of phase estimation is counterproductive, leading to a reduced quantum advantage at small values of $\mathcal{D}$. In fact, at small $\mathcal{D}$ the collection efficiency becomes $\eta_{c}(\mathcal{D})\ll 1$, and $k^{(\tau)}_{\text{opt}}(\mathcal{D})\ll 1$, resulting in quantum performance comparable to classical measurements. However, in amplitude SSNQI this choice still provides the optimal noise reduction, although modest, at small spatial scale  Ruo-Berchera et al. (2020); Ortolano et al. (2023b).
@@
In contrast, since high-frequency fluctuations are inherently suppressed in NIQPI through the solution of the TIE, it is more convenient to apply the factor $k^{(\text{TIE})}_{\text{opt}}\approx\eta_{0}$, which more effectively optimizes noise suppression at low spatial frequencies, even to high-resolution intensity images.
The quantum advantage is most pronounced at small defocus distances $dz$, where the intensity variations induced by phase propagation are weaker. In this regime, quantum noise suppression becomes particularly impactful, achieving an advantage close to 30%. We recall that, in the opposite regime of large refocusing distance, the finite-difference approximation is poor, thereby reducing both the resolution and the quantitative accuracy of phase estimation Ortolano et al. (2023b).
@@
In Fig. 2, we present the classical and quantum reconstructions of both the phase and transmittance profiles. The reconstructions clearly demonstrate the effectiveness of the method in separating transmittance and phase information, as well as highlighting the quantum advantage. For the transmittance reconstruction a resolution parameter $\mathcal{D}=3.8$, corresponding to a spatial resolution of $18\,\mu\text{m}$, was used to maximize the collection efficiency $\eta_{c}(\mathcal{D})$. This choice enabled the achievement of an NRF of 0.45 and a quantum advantage of 25% in the standard deviation of the reconstructed amplitude. Note that, given the transmittance level of the mask, $\tau=0.94$, and the large photon counts, the classical image already provides good retrieval, which limits the visual perception of the quantum advantage in the reconstructed image. However, this does not diminish the significant quantum advantage observed in the measured uncertainty over the transmittance.
@@
Conversely, the phase reconstruction reveals a visual quantum advantage. The phase’s values estimated for the smallest defocus distance $dz$, where the approximation in Eq. (3) is most accurate, are listed in the caption. They show a good agreement with theoretical predictions and a reduced uncertainty for the quantum case.
Fig. 3 presents a quantitative evaluation of the spatial phase resolution achieved with the new $k^{(\text{TIE})}_{\text{opt}}$-based phase retrieval strategy, as a function of the effective resolution of the intensity images, $\mathcal{D}$. The results, shown in Fig. 3a, are reported for different defocus distances $dz$. Each data point represents the experimental resolution extracted from phase images averaged over $10^{3}$ frames to eliminate residual shot-noise. The shaded area indicates the theoretical resolution limit achievable with the current optical system. Fig. 3b displays representative phase reconstructions, also averaged over $10^{3}$ frames, for two defocus distances, $dz=0.025\,\text{mm}$ and $dz=0.1\,\text{mm}$, at three different values of $\mathcal{D}$. Rectangular regions within each phase image are magnified to illustrate resolution variations more clearly.
To quantify the spatial resolution, horizontal phase profiles were extracted along the dashed lines indicated in Fig. 3b. These profiles were then fitted using an error function (Edge Spread Function, ESF), allowing the spatial resolution to be defined as $r_{\text{phase}}=2\sqrt{2\ln(2)}\,w$, where $w$ denotes the characteristic width parameter obtained from the ESF fit (see M&M for further details). The uncertainty associated with the resolution measurements was estimated from the 95% confidence interval of the parameter $w$. As clearly shown by the curves in Fig. 3a, decreasing the defocus distance $dz$ improves the resolution by enabling a more accurate approximation of the derivative term in Eq. (2), ultimately achieving a phase spatial resolution of approximately $4\,\mu$m for $dz=0.0125$ mm and for the minimal integration lenght $L_{\text{det}}$ of the physical pixel, that correspond to $\mathcal{D}=0.32$. Conversely, increasing the integration area leads to a loss of resolution, where all the curves approach the same asymptotic behavior, with a spatial resolution of the phase images around $18\,\mu\text{m}$, as achieved in our previous work Ortolano et al. (2023b). As mentioned this strongly limited practical applications of this technique. In the present work, this limitation was overcome by fully exploiting the unique properties of TIE noise propagation at high spatial frequencies.
@@
Figure 4: Biological quantum phase imaging with high resolution. (Left column) Phase images obtained by averaging 100 frames, serving as a shot-noise-free reference. (Middle column) Classical single-shot phase images, affected by shot-noise. (Right column) Quantum-corrected single-shot phase images, showing visibly improved quality and enhanced contrast compared to the classical case. Each row corresponds to a different biological sample, and the same color scale is used within each row. The scale bar corresponds to $20~{}\mu\text{m}$.
@@
Fig.  4 presents wide-field phase reconstructions of ova at different developmental stages: from unfertilized eggs (top row), to fertilized eggs in early cleavage stages (middle row), to advanced embryonic forms (bottom row). These three biological conditions were imaged under different experimental settings. The defocus distances used in the TIE reconstruction were $dz=10\,\mu\text{m}$, $5\,\mu\text{m}$, and $5\,\mu\text{m}$, respectively, and the average number of detected photons per pixel was approximately 1000, 300, and 600 for the three samples.
The first column displays the reference phase image obtained by averaging 100 independent frames in order to eliminate any shot-noise. The second column shows the phase reconstruction from a single frame without any quantum correction. The third column presents the quantum-enhanced phase image obtained by the substitution in Eq. (4) using the optimized correction factor $k^{(\text{TIE})}_{\text{opt}}=\eta_{0}$.
The quantum-corrected images demonstrate a visible improvement in contrast and noise suppression compared to the classical single-shot reconstructions, approaching the quality of the 100-frame averaged images. In particular, a quantum advantage between 5% and 20% was measured across different biological samples. These results highlight both the robustness of the method in the presence of low signal levels and its capacity to provide high-fidelity phase reconstructions in biologically relevant conditions.
@@ II Conclusion
We have demonstrated a powerful sub-shot-noise technique in the domain of phase imaging, which does not suffer from any trade-off between spatial resolution and quantum gain in sensitivity. This enables the development of a microscope with diffraction-limited resolution and sub-shot-noise enhancement. In the microscopic system presented here, although we reach micrometer-scale resolution, the minimal achievable resolution does not yet match the diffraction limit of the imaging system. This discrepancy is due to technical limitations, the most relevant being the imperfect approximation of the derivative in Eq. (3) for the chosen value of $dz$. Further improvements could be obtained by multi-distance approaches to TIE, which combine intensity measurements at both small and large defocus distances Zuo et al. (2013), by employing structured illumination techniques Zuo et al. (2017) or simply by considering a smaller defocusing distance. This last approach would be possible just by a finer control of the translator stages. Additionally, in our setup, the diffraction limit is close to the physical pixel size of the detector, making it challenging to precisely evaluate the resolution near the theoretical limit.
To demonstrate the applicability of our quantum-enhanced phase imaging method to real-world biological systems, we performed phase imaging of cells with details of a few micrometers, clearly showing less noise artifacts with respect to the classical measurement. Finally, our method is scanning free, requires only two shots for complete phase retrieval and do not need complicated interferometric setups. Filling the gap between proof-of-concept and real word applicability, this work represents an important step for the quantum sensing technology, where only a very limited number of protocols have reached an application level like ours. Several applications can already be envisaged whenever high illumination level can damage the sample or in fast imaging where the illumination level for the single frame should remain limited.
@@ III.1 Experimental layouts
The experimental setup is based on a continuous-wave (CW) laser source (OBIS405 Coherent) operating at a wavelength $\lambda_{p}=405\,\text{nm}$ with a beam width of $w=0.5\,\text{mm}$. Although the source is CW, it is triggered in a pulsed mode synchronized with the camera’s acquisition, with a maximum output power of 100 mW. The pump beam generates photon pairs via spontaneous parametric down-conversion (SPDC) in a 1.5 cm-long, type-II beta-barium borate (BBO) nonlinear crystal with transverse dimensions of 0.8 $\times$ 0.8 cm2. The down-converted photons are filtered by an interferential filter centered at $800\pm 20$ nm, selecting only photons near the degenerate wavelength $\lambda_{d}=810\,\text{nm}$.
Spatial correlations are established at the crystal far-field (CFF) plane, formed by a lens with focal length $f=1\,\text{cm}$. A second imaging lens with focal length $f=1.6\,\text{cm}$ projects the far-field onto the CCD detection plane with a magnification factor of $M=8$. The detector is a Pixis 400BR Excelon CCD camera (Princeton Instruments) with a $1024\times 1024$ pixel array and $13\,\mu\text{m}$ pixel pitch. The pixel size on the object plane is approximately $1.5\,\mu\text{m}$. The camera operates in linear mode with high quantum efficiency ($>95\%$), low electronic noise, and full fill factor.

### v2

@@ I Introduction
To understand why this is the case, it is essential to consider how intensity noise propagates into the retrieved phase profile. By examining Eq. (2), under the assumption that the finite-difference term $I(\mathbf{x},dz)-I(\mathbf{x},-dz)$ is dominated by noise fluctuations $\sigma(\mathbf{x})$, and considering a nearly uniform illumination intensity $I(\mathbf{x},0)\approx I_{0}$, the equation simplifies to [46]:
$k\frac{\sigma(\mathbf{x})}{\sqrt{2}\,I_{0}\,dz}=\nabla^{2}\phi_{\text{noise}}(\mathbf{x}).$ (5)
@@
where $\tilde{\sigma}(\mathbf{q})$ and $\tilde{\phi}_{\text{noise}}(\mathbf{q})$ denote the Fourier transforms of the intensity noise and the resulting phase noise, respectively, and $\mathbf{q}$ represents the spatial frequency vector. In the case of shot-noise we can assume a flat spectrum, $\tilde{\sigma}(\mathbf{q})=\tilde{\sigma}$, for $\mathbf{q}\neq 0$. The presence of the $|\mathbf{q}|^{2}$ term in the denominator highlights that the phase noise in the TIE reconstruction is dominated by low spatial-frequency components, i.e. by the slowly varying shot-noise fluctuations across the image. In contrast, high-frequency noise components, such as pixel-to-pixel shot-noise, are strongly suppressed by the intrinsic low-pass filtering nature of the TIE, and therefore have minimal impact on the retrieved phase. Incidentally, the remaining low-frequency components of the shot-noise are precisely those efficiently removed through quantum correlations, by the substitution in Eq. (4).
As a direct consequence of the TIE noise propagation described in Eq. (6), it is not necessary to intentionally reduce the spatial resolution of the detected intensity images prior to applying the TIE by increasing $L_{\mathrm{det}}$, in order to achieve an effective quantum noise reduction. At the same time, it is worth noting that the intrinsic suppression of high-frequency noise does not prevent the retrieval of high-resolution phase images. In fact, as it is clear from Eq. (2), in the direct-propagation problem, higher-frequency components of the phase object produce a stronger effect on the intensity. Thus, regions with rapid phase variations, i.e. high-frequency features, are usually reconstructed accurately despite the suppression introduced by solving the TIE [46].
@@
These results are shown in Fig. 1c. The four panels analyze the quantum advantage as a function of the phase resolution for four different defocusing distances $dz$. The quantum advantage is quantified as the ratio between two Pearson correlation coefficients: one calculated using quantum-corrected phase images, $\phi^{\text{qua}}(\mathbf{x})$ and the other using classical (shot-noise-limited) phase images, $\phi^{\text{cla}}(\mathbf{x})$, each compared with a reference, $\phi_{r}(\mathbf{x})$, obtained by averaging $10^{3}$ frames to suppress residual shot-noise fluctuations. The classical and quantum averaged images, obtained by sending a total number of photons through the sample that is one thousand times larger than in the single-shot case, are indistinguishable for any practical purpose. From a numerical standpoint, either image can therefore be used equivalently as a noise-free reference. The Pearson correlation coefficient $\mathcal{C}$, a widely used metric to evaluate the similarity between two images, is defined as:
$\mathcal{C}=\frac{\sum_{\mathbf{x}}\left(\phi_{r}(\mathbf{x})-\overline{\phi}_{r}\right)\left(\phi(\mathbf{x})-\overline{\phi}\right)}{\sqrt{\text{Var}[\phi_{r}]\,\text{Var}[\phi]}},$ (7)
@@
In the figure, the blue data points represent the experimental quantum advantage obtained using the optimization factor $k^{(\text{TIE})}_{\text{opt}}=\eta_{0}$ in Eq. (4), with a single-channel overall detection efficiency estimated as $\eta_{0}\approx 0.7$, while the red points indicate results obtained with $k^{(\tau)}_{\text{opt}}=\eta_{0}\eta_{c}(\mathcal{D})$, used for transmittance measurements and previously employed for phase imaging in Ref. [45]. The shaded red and blue areas represent results from Fourier-optics simulations [67], showing good agreement with the experimental data, whereas the green area indicates the gain in the quantum advantage achievable with the new reconstruction strategy.
For each defocusing distance $dz$, the data are analyzed as a function of the phase spatial resolution. Since the phase resolution is directly determined by the spatial resolution of the intensity measurements, the latter is simply varied by increasing the effective integration area $L_{\mathrm{det}}$ (i.e., $\mathcal{D}$). This is implemented by applying an averaging filter over neighborhoods ranging from 1 to 12 pixels. The phase resolution is evaluated by fitting horizontal phase profiles, obtained from phase images averaged over $10^{3}$ frames, with an error function (Edge Spread Function, ESF). The spatial resolution is then defined as $r_{\mathrm{phase}}=2\sqrt{2\ln 2}\,w,$ where $w$ is the characteristic width parameter obtained from the ESF fit (see M&M for better details). The averaging procedure over different frames is used to suppress shot-noise and isolate the intrinsic phase resolution. This choice allows us to disentangle the intrinsic spatial resolution of the optical system and of the TIE reconstruction from noise-induced effects.
For all values of $dz$, increasing $\mathcal{D}$ degrades the phase spatial resolution. At the same time considering smaller distances $dz$, the finite-difference approximation of the axial derivative in Eq. (2) is more accurate, enabling a higher spatial resolution (see M&M for details). In the opposite regime of large defocus distances, the finite-difference approximation deteriorates, resulting in a simultaneous reduction of both spatial resolution and quantitative accuracy of the phase estimation [45].
@@
We observe that using the correction factor $k^{(\tau)}_{\mathrm{opt}}(\mathcal{D})$, originally introduced for amplitude imaging, instead of $k^{(\text{TIE})}_{\text{opt}}$, is counterproductive in the context of phase estimation. In particular, at the highest phase resolution, corresponding to small integration area $\mathcal{D}$, the collection efficiency is $\eta_{c}(\mathcal{D})\ll 1$, leading to $k^{(\tau)}_{\mathrm{opt}}(\mathcal{D})\ll 1$. Consequently the quantum performance is comparable to the classical case, according to Eq. (4). While this choice remains optimal for amplitude SSNQI, where it still yields the best noise reduction  [54, 45], it is not suited for phase reconstruction.
@@
In contrast, in NIQPI, high-spatial-frequency fluctuations are intrinsically suppressed by the solution of the TIE. This makes it more advantageous to employ the factor $k^{(\mathrm{TIE})}_{\mathrm{opt}}\approx\eta_{0}$, which preferentially optimizes noise suppression at low spatial frequencies, even when operating on high-resolution intensity images.
[…]
@@
In Fig. 2, we present the classical and quantum reconstructions of both the phase and transmittance profiles. The reconstructions clearly demonstrate the effectiveness of the method in separating transmittance and phase information, as well as highlighting the quantum advantage. For the transmittance reconstruction a $\mathcal{D}=3.8$ obtained by averaging over 12 pixels was used in order to maximize the collection efficiency $\eta_{c}(\mathcal{D})$. This choice enabled the achievement of an NRF of 0.45 and a quantum advantage of 25% in the standard deviation of the reconstructed amplitude. Note that, given the transmittance level of the mask, $\tau=0.93$, and the large photon counts, the classical image already provides good retrieval, which limits the visual perception of the quantum advantage in the reconstructed image. However, this does not diminish the significant quantum advantage observed in the measured uncertainty over the transmittance.
@@
Conversely, the phase reconstruction are retrieved without any integration on the detector’s area and with the new optimized correction factor $k^{(\mathrm{TIE})}_{\mathrm{opt}}$, revealing a visual quantum advantage. The phase’s values estimated for the smallest defocus distance $dz$, are listed in the caption. They show a good agreement with theoretical predictions and a reduced uncertainty for the quantum case. For higher $dz$ the approximation in Eq. (3) is less accurate, and other technical issues in the alignment affect the phase estimation, as discusses in the Supplemental Document.
@@
Figure 3: Biological quantum phase imaging (a) (Left column) Phase images obtained by averaging 100 frames, serving as a shot-noise-free reference. (Middle column) Classical single-shot phase images, affected by shot-noise. (Right column) Quantum-corrected single-shot phase images showing visibly improved quality and enhanced contrast compared to the classical case. Each row corresponds to a different biological sample, and the same color scale is used within each row. (b) Two areas, corresponding to the square boxes in last row of (a), are cropped. The color scales represent the entire range of the retrieved phase values. Green and blue dots are specific positions chosen for the quantitative analysis. (c) (Left) The average phase values corresponding to the six dots calculated over 100 frames. (Right) Standard deviation of the phase values in correspondence of the six dots.
@@
Fig. 3a presents wide-field phase reconstructions of ova at different developmental stages: from unfertilized eggs (top row), to fertilized eggs in early cleavage stages (middle row), to advanced embryonic forms (bottom row). These three biological conditions were imaged under different experimental settings. The defocus distances used in the TIE reconstruction were $dz=10\,\mu\text{m}$, $5\,\mu\text{m}$, and $5\,\mu\text{m}$, respectively, and the average number of detected photons per pixel was approximately 1000, 300, and 600 for the three samples.
The first column displays the reference phase image obtained by averaging 100 independent frames in order to eliminate any shot-noise. The second column shows the phase reconstruction from a single frame without any quantum correction. The effect of shot-noise in the phase reconstruction manifests as a low frequency cloud like patterns producing under- and over-exposed regions. The third column presents the quantum-enhanced phase image obtained by subtracting the shot-noise measured in the reference beam according to Eq. (4) with $k^{(\text{TIE})}_{\text{opt}}=\eta_{0}$.
The quantum-corrected images demonstrate a visible improvement in contrast and noise suppression compared to the classical single-shot reconstructions, approaching the quality of the 100-frame averaged images. Two small portions of the last image, the ones inside the square boxes, are zoomed in Fig 3b. Even in these small details, and using a natural color scale, the improved fidelity of the quantum image is clearly appreciable: several details, lost passing from the 100 frames to the single classical frame, are restored in the quantum frame. Moreover, we have chosen six points in the image (green and blue dots) and reported the corresponding average phase values and their standard deviation in Fig. 3c. While the average values are the same for the classical and quantum cases, the standard deviation is clearly lower for the quantum reconstructions, showing a quantum advantage in the range 20-25% in the quantitative estimation of the phase point-by-point. Concerning the overall fidelity, which includes the recovered morphological information and is well captured by the Pearson correlation coefficient $\mathcal{C}$, a quantum advantage between $5\%$ and $20\%$ is observed across the different biological samples. The variation in the measured advantage mainly reflects differences in sample absorption and thickness, which lead to different signal levels and thus to a varying impact of shot-noise. These results highlight both the robustness of the method in the presence of low signal levels and its capacity to provide high-fidelity phase reconstructions in biologically relevant conditions.
@@ II Conclusion
We have demonstrated a powerful sub-shot-noise technique in the domain of phase imaging, which does not suffer from any trade-off between spatial resolution and quantum gain in sensitivity. This enables the development of a microscope with diffraction-limited resolution and sub-shot-noise enhancement. In the microscopic system presented here, although we reach micrometer-scale resolution, the minimal achievable resolution does not yet match the diffraction limit of the imaging system. This discrepancy is due to technical limitations, the most relevant being the imperfect approximation of the derivative in Eq. (3) for the chosen value of $dz$. Further improvements could be obtained by multi-distance approaches to TIE, which combine intensity measurements at both small and large defocus distances [71], by employing structured illumination techniques [73] or simply by considering a smaller defocusing distance. This last approach would be possible just by a finer control of the translator stages. Additionally, in our setup, the diffraction limit is close to the physical pixel size of the detector, making it challenging to precisely evaluate the resolution near the theoretical limit.
To demonstrate the applicability of our quantum-enhanced phase imaging method to real-world biological systems, we performed phase imaging of cells with details of a few micrometers, clearly showing less noise artifacts with respect to the classical measurement. We achieved a quantum advantage of up to 20 % in terms of image fidelity, evaluate through the correlation coefficient, and up to 25% in the quantitative phase uncertainty. Finally, our method is scanning free, requires only two shots for complete phase retrieval and do not need complicated interferometric setups. Filling the gap between proof-of-concept and real word applicability, this work represents an important step for the quantum sensing technology, where only a very limited number of protocols have reached an application level like ours. Several applications can already be envisaged whenever high illumination level can damage the sample or in fast imaging where the illumination level for the single frame should remain limited.
@@ III.1 Experimental layouts
The experimental setup is based on a continuous-wave (CW) laser source (OBIS405 Coherent) operating at a wavelength $\lambda_{p}=405\,\text{nm}$ with a beam width of $0.5\,\text{mm}$. Although the source is CW, it is triggered in a pulsed mode synchronized with the camera’s acquisition, with a maximum output power of 100 mW. The pump beam generates photon pairs via spontaneous parametric down-conversion (SPDC) in a 1.5 cm-long, type-II beta-barium borate (BBO) nonlinear crystal with transverse dimensions of 0.8 $\times$ 0.8 cm2. The down-converted photons are filtered by an interferential filter centered at $800\pm 20$ nm, selecting only photons near the degenerate wavelength $\lambda_{d}=810\,\text{nm}$.
Spatial correlations are established at the crystal far-field (CFF) plane, formed by a lens with focal length $f=1\,\text{cm}$. A second imaging lens with focal length $f=1.6\,\text{cm}$ projects the far-field onto the CCD detection plane with a magnification factor of $M=8$. The detector is a Pixis 400BR Excelon CCD camera (Princeton Instruments) with a $1024\times 1024$ pixel array and $13\,\mu\text{m}$ pixel pitch. The pixel size on the object plane is approximately $1.5\,\mu\text{m}$. The camera operates in linear mode with high quantum efficiency ($95\%$), low electronic noise, and full fill factor.